# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 4.2 Walls in the space of constraints
(2i) Let $\widehat{C}\simeq{\mathbb{P}}^{1}$ and $\widehat{\boldsymbol{w}}(t^{*})$ consist of $n+1$ distinct points. If all points $w_{i}^{\sigma}(t^{*})$ are smooth, then we get the case (2i).
@@
Suppose that $C={\Phi}({\mathbb{P}}^{1})$ is singular at some point $w_{i}^{\sigma}(t^{*})$. Note that $\xi$ must be unibranch at each point of $\widehat{\boldsymbol{w}}(t^{*})$. Since $\xi$ is a genetic element of an $(n-1)$-dimensional family in ${\mathcal{M}}_{0,n}(\Delta)$, fixing the position of $w_{i}^{\sigma}$, we obtain a family of dimension $\geq n-2$. Applying [13, Inequality (5) in Lemma 2.1], we obtain
$\sum_{\sigma\subset\partial P}\|\sigma\|_{\mathbb{Z}}\geq 2+\left(\sum_{\sigma\subset\partial P}\|\sigma\|_{\mathbb{Z}}-n\right)+(n-3)+\sum_{B}({\operatorname{ord}}B-1)$
$=\sum_{\sigma\subset\partial P}\|\sigma\|_{\mathbb{Z}}-1+\sum_{B}({\operatorname{ord}}B-1)\ ,$
where $B$ ranges over all singular branches of $\xi$, and hence $\xi$ has a unique singular branch, and this branch is centered at $w_{i}^{\sigma}(t^{*})$ and has multiplicity $2$. Thus, the type of the singularity must be $A_{2m}$, $m\geq k_{i}^{\sigma}$.
@@
(2ii) Suppose that $\widehat{C}\simeq{\mathbb{P}}^{1}$ and some of the points of the sequence $\widehat{\boldsymbol{w}}(t^{*})$ coincide. For the dimension reason, the number #$\#(\widehat{\boldsymbol{w}}(t^{*}))$ of points in $\widehat{\boldsymbol{w}}(t^{*})$ is equal to $n-1$ (i.e., $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})=w$ for some $\sigma\subset\partial P$ and $i\neq j$), and all these $n-1\geq 2$ points are in general position subject to the Menelaus relation (1). If $n-1=2$, then the claim of item (2ii) is fulfilled by [22, Lemma 3.5]. Assume that $n-1\geq 3$. If $C={\Phi}(\widehat{C})$ is unibranch at each point of $\widehat{\boldsymbol{w}}(t^{*})$, then, by Lemma 2.4, the curve $C$ is immersed and smooth along the toric divisors. If $C$ is not unibranch at $w$, then by Lemma 2.4 it has two local branches at $w$; furthermore, $C$ must be unibranch at each point $w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\neq w$ and, in addition, smooth if $k_{i^{\prime}}^{\sigma^{\prime}}\geq 2$ due to Lemma 2.4 and claim (1ii). Let us show that $C$ is immersed. Fixing the position of $w$ and the position of one more point $w^{\prime}\in{\boldsymbol{w}}(t^{*})\setminus\{w\}$, we obtain a family of dimension at least $n-3\geq 1$; thus, [13, Inequality (5) in Lemma 2.1] applies:
$c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})\geq 2+(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-n+2)+\sum_{B}({\operatorname{ord}}B-1)+(n-4)$
$=c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+\sum_{B}({\operatorname{ord}}B-1)\ ,$
@@
where $B$ runs over all singular local branches of $C$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w,w^{\prime}\}$, and hence $C$ is immersed.
@@
(2iii) We are left with the case of $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq\widehat{C}_{2}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{1}\cap\widehat{C}_{2}=\{p\}$ is one point. For the dimension reason, the points of $\widehat{\boldsymbol{w}}(t^{*})$ are in general position subject to exactly two Menelaus conditions (induced by the components of $\widehat{C}$), and they are all distinct. Each of the curves $C_{1}$, $C_{2}$, which passes through at least three points of $\widehat{\boldsymbol{w}}(t^{*})$, is immersed and smooth along the toric divisors and has even intersection multiplicity with ${\operatorname{Tor}}(\partial P_{\Delta})$ at any point, by Lemma 2.4. If $C_{1}$ or $C_{2}$ passes through exactly two points of $\widehat{\boldsymbol{w}}(t^{*})$, then it is a multiple covering of a line as described in item (2iv). Let ${\Phi}(p)\in{\operatorname{Tor}}(P_{\Delta})^{\times}$. Then, by Lemma 2.4, the curves $C_{1}$ and $C_{2}$ do not share points in $\widehat{\boldsymbol{w}}(t^{*})$, and we claim that all their intersection points are ordinary nodes. Due to the genericity assumptions for $\widehat{\boldsymbol{w}}(t^{*})$, we have to study the only case of both $C_{1}$ and $C_{2}$ immersed. If we freely move the points of $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})$ so that the corresponding Menelaus condition induced by $C_{1}$ retains, and fix the curve $C_{2}$, then the persisting tangency condition of a germ ${\Phi}:(\widehat{C}_{1},q)\to{\operatorname{Tor}}(P_{\Delta})$ to the curve $C_{2}$ would yield that the tangent space to the considered family of curves in the linear system $|C_{1}|$ would be contained in $H^{0}(\widehat{C}_{1},{\mathcal{O}}_{\widehat{C}_{1}}({\boldsymbol{d}}_{1}))$, where
$\deg{\boldsymbol{d}}_{1}=C_{1}^{2}-(C_{1}^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{1}]+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{1}]-\#({\boldsymbol{w}}(t^{*})\cap C_{1}))-1$
@@
(2iv) Suppose that ${\Phi}:{\mathbb{P}}^{1}\to{\operatorname{Tor}}(P_{\Delta})$ is an $s$-multiple covering onto its image, $s\geq 2$. For the dimension reason, the sequence $\widehat{\boldsymbol{w}}(t^{*})$ contains at least $n$ distinct points. If $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n-1$ distinct points, then at each point of $\widehat{\boldsymbol{w}}(t^{*})$ we have the ramification index $s$; hence, by the Riemann-Hurwitz formula, we get
@@ 4.3 Wall-crossing
Let $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ be a generic path in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$, and let $t^{*}\in(0,1)$ be such that ¯$\overline{\mathcal{M}}_{0,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ contains an element $\xi$ as described in one of the items of Lemma 4.2(2). Then, for each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the numbers $W_{0}^{\kappa}(t)=W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ do not change as $t$ varies in a neighborhood of $t^{*}$.
@@
(1) Suppose that $\xi\in\overline{\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ is as in Lemma 4.2(2i) and, moreover, if it has a singular branch at $w_{i}^{\sigma}(t^{*})$, then $k_{i}^{\sigma}=1$. We derive the constancy of $W_{0}^{\kappa}(t)$, $|t-t^{*}|<\delta$ from [12, Lemma 15] and [13, Lemma 2.4(1)]. To this end, we have to establish the following transversality statement (cf. [12, Lemma 13]). Choose a sufficiently large integer $s$. For each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$, we set $I_{z}=I^{cond}(C,z)/{\mathfrak{m}}_{z}^{s}\subset{\mathcal{O}}_{C,z}/{\mathfrak{m}}_{z}^{s}$, the quotient of the conductor ideal by the power of the maximal ideal, which can be viewed as the tangent cone to the stratum parameterizing equigeneric deformations (see [3, Theorem 4.15]). For each point $w=w_{i}^{\sigma}(t^{*})$, where $C$ is smooth and which is fixed, respectively, mobile, we set
$I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ (\varphi\cdot C)_{w}\geq 2k_{i}^{\sigma}\}/{\mathfrak{m}}_{w}^{s}\subset{\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}\ ,$
@@
which can be viewed as the tangent space to the stratum parameterizing deformations that keep the intersection number with ${\operatorname{Tor}}(\sigma)$ at $w$, respectively, in a nearby point on ${\operatorname{Tor}}(\sigma)$. For each point $w=w_{i}^{\sigma}(t^{*})$ with $k_{i}^{\sigma}=1$, where $C$ is singular (i.e., has a singular branch of type $A_{2m}$, $m\geq 1$) and which is fixed, respectively, mobile, we set
$I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ (\varphi\cdot C)_{w}\geq 2+2m\}/{\mathfrak{m}}_{w}^{s}\subset{\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}\ ,$

### v2

@@ 4.2 Walls in the space of constraints
(2i) Let $\widehat{C}\simeq{\mathbb{P}}^{1}$, $\widehat{\boldsymbol{w}}(t^{*})$ consist of $n+1$ distinct points, and let ${\Phi}$ take $\widehat{C}$ birationally onto $C={\Phi}(\widehat{C})\subset{\operatorname{Tor}}(P_{\Delta})$.
First, we consider the minimal value $n=2$. In such a case, $P_{\Delta}$ is a triangle and, for each $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$, there exists a unique curve $C={\Phi}({\mathbb{P}}^{1})$ (where $\xi=[{\Phi}:({\mathbb{P}}^{1},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ which is smooth along the toric divisors and is nodal, while the number of elliptic nodes in ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ does not depend on the choice of $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$, see [28, Lemma 3.5].
@@
Suppose that $n\geq 3$. Note that $\xi$ must be unibranch at each point of $\widehat{\boldsymbol{w}}(t^{*})$. Since $\xi$ is a generic element of an $(n-1)$-dimensional family in ${\mathcal{M}}_{0,n}(\Delta)$, fixing the position of a point $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}(t^{*})$, we obtain a family of dimension $\geq n-2\geq 1$. Applying [16, Inequality (5) in Lemma 2.1], we get
$-CK_{{\operatorname{Tor}}(P_{\Delta})}\geq 2+\left(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n\right)+(n-3)+\sum_{B}({\operatorname{mult}}Q-1)$
$=-CK_{{\operatorname{Tor}}(P_{\Delta})}-1+\sum_{B}({\operatorname{mult}}Q-1)\quad\Longrightarrow\quad\sum_{B}({\operatorname{mult}}Q-1)\leq 1\ ,$ (13)
where $B$ is the set of all singular branches of $\xi$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w_{i}^{\sigma}\}$. In particular, if $C$ is singular at some point $w_{i}^{\sigma}$, the singularity is of type $A_{2m}$. The requirements of item (2i) are fulfilled.
@@
(2ii) Suppose that $\widehat{C}\simeq{\mathbb{P}}^{1}$ is birationally taken by ${\Phi}$ onto its image in ${\operatorname{Tor}}(P_{\Delta})$, and some of the points of the sequence $\widehat{\boldsymbol{w}}(t^{*})$ coincide. The latter is possible only if $n\geq 3$. For the dimension reason, the number #$\#(\widehat{\boldsymbol{w}}(t^{*}))$ of points in $\widehat{\boldsymbol{w}}(t^{*})$ is equal to $n$ (i.e., $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})=w$ for some $\sigma\subset\partial P_{\Delta}$ and $i\neq j$), and all these $n\geq 3$ points are in general position on ${\operatorname{Tor}}(\partial P_{\Delta})$ subject to the Menelaus relation (1). If $C={\Phi}(\widehat{C})$ is unibranch at each point of $\widehat{\boldsymbol{w}}(t^{*})$, then, by Lemma 2.4, the curve $C$ is immersed and smooth along the toric divisors. If $C$ is not unibranch at $w$, then by Lemma 2.4 it has two local branches at $w$; furthermore, $C$ must be unibranch at each point $w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\neq w$. So, assume that $C$ has two local branches at $w$.
If $n-1=2$, then $P_{\Delta}$ is a lattice triangle. Using affine automorphisms of ${\mathbb{Z}}^{2}$, we can take $P_{\Delta}={\operatorname{conv}}\{(m_{1},0),(m_{1}+k_{1}+k_{2},0),(0,m_{2})\}$ with $m_{1},m_{2},k_{1},k_{2}$ even and write a parameterization of $C$ in the form
$x=a\tau^{m_{2}},\quad y=b\tau^{m_{1}}(\tau-1)^{k_{1}}(\tau-\theta)^{k_{2}},\quad a,b,\theta\in{\mathbb{R}}^{*}.$
Since $x(1)=x(\theta)$, we obtain $\theta=-1$. Note that $k_{1}\neq k_{2}$, since otherwise, $x$ and $y$ would be functions of $\tau^{2}$, that is, ${\Phi}:\widehat{C}\to C$ would be a double covering against the initial assumption. Relations
$x=a(1+m_{2}(\tau-1)+O((\tau-1)^{2}),\quad x=a(1-m_{2}(\tau+1)+O((\tau+1)^{2}),$
show that the local branches of $C$ at $w$ are smooth. Furthermore,
$x=a\tau^{m_{2}},\quad y=b\tau^{m_{1}}+b(k_{2}-k_{1})\tau^{m_{1}+1}+O(\tau^{m_{1}+2});$
hence, $C$ is smooth at the two remaining points of $\widehat{\boldsymbol{w}}$. In addition, $C$ is immersed everywhere, since $\frac{dx}{d\tau}\neq 0$ for $\tau\neq 0$.
Let $n-1\geq 3$. We show that $C$ is immersed. Fixing the position of $w$ and the position of one more point $w^{\prime}\in{\boldsymbol{w}}(t^{*})\setminus\{w\}$, we obtain a family of dimension at least $n-3\geq 1$; thus, [16, Inequality (5) in Lemma 2.1] applies:
$-CK_{{\operatorname{Tor}}(P_{\Delta})}\geq 2+(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+2)+\sum_{B}({\operatorname{mult}}Q-1)+(n-4)$
$=-CK_{{\operatorname{Tor}}(P_{\Delta})}+\sum_{B}({\operatorname{mult}}Q-1)\ ,$
@@
where $B$ is the set of all singular local branches of $C$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w,w^{\prime}\}$, and hence $C$ is immersed.
@@
(2iii) Consider the case of $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq\widehat{C}_{2}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{1}\cap\widehat{C}_{2}=\{p\}$ is one point. For the dimension reason, the points of $\widehat{\boldsymbol{w}}(t^{*})$ are in general position subject to exactly two Menelaus conditions (induced by the components of $\widehat{C}$), and they are all distinct. Each of the curves $C_{1}$, $C_{2}$, which passes through at least three points of $\widehat{\boldsymbol{w}}(t^{*})$, is immersed and smooth along the toric divisors and has even intersection multiplicity with ${\operatorname{Tor}}(\partial P_{\Delta})$ at any point, by Lemma 2.4. If $C_{1}$ or $C_{2}$ passes through exactly two points of $\widehat{\boldsymbol{w}}(t^{*})$, then it is a multiple covering of a line as described in item (2iv). Let ${\Phi}(p)\in{\operatorname{Tor}}(P_{\Delta})^{\times}$. Then, by Lemma 2.4, the curves $C_{1}$ and $C_{2}$ do not share points in $\widehat{\boldsymbol{w}}(t^{*})$, and we claim that all their intersection points are ordinary nodes. Due to the genericity assumptions for $\widehat{\boldsymbol{w}}(t^{*})$, we have to study the only case of both $C_{1}$ and $C_{2}$ immersed. If we freely move the points of $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})$ so that the corresponding Menelaus condition induced by $C_{1}$ retains, and fix the curve $C_{2}$, then the persisting tangency condition of a germ ${\Phi}:(\widehat{C}_{1},q)\to{\operatorname{Tor}}(P_{\Delta})$ to the curve $C_{2}$ would yield that the tangent space to the considered family of curves in the linear system $|C_{1}|$ would be contained in $H^{0}(\widehat{C}_{1},{\mathcal{O}}_{\widehat{C}_{1}}({\boldsymbol{d}}_{1}))$, where
$\deg{\boldsymbol{d}}_{1}=C_{1}^{2}-(C_{1}^{2}+C_{1}K_{{\operatorname{Tor}}(P_{\Delta})}+2)-(-C_{1}K_{{\operatorname{Tor}}(P_{\Delta})}-\#({\boldsymbol{w}}(t^{*})\cap C_{1}))-1$
@@
(2iv) Suppose that ${\Phi}:{\mathbb{P}}^{1}\to{\operatorname{Tor}}(P_{\Delta})$ is an $s$-multiple covering onto its image, $s\geq 2$. For a generic $t\in[0,1]$, $\widehat{\boldsymbol{w}}(t)$ consists of $n+1$ distinct points; hence the sequence $\widehat{\boldsymbol{w}}(t^{*})$ contains at least $n$ distinct points. If $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n+1$ distinct points, then at each point of $\widehat{\boldsymbol{w}}(t^{*})$ we have the ramification index $s$; thus, by the Riemann-Hurwitz formula, we get
@@ 4.3 Wall-crossing
Let $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ be a generic path in $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$, and let $t^{*}\in(0,1)$ be such that ¯$\overline{\mathcal{M}}_{0,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ contains an element $\xi$ as described in one of the items of Lemma 4.2(2). Then, for each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the numbers $W_{0}^{\kappa}(t)=W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ and $\widetilde{W}_{0}^{\kappa}(t)=\widetilde{W}_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ do not change as $t$ varies in a neighborhood of $t^{*}$.
@@
The proof is separated into parts (1)-(6) in accordance with the classification of walls in Lemma 4.2(2) and specific degenerations occurring on these walls.
@@
(1) Suppose that $\xi\in\overline{\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ is as in Lemma 4.2(2i), and $\xi$ has a singular local branch centered in ${\operatorname{Tor}}(P_{\Delta})^{\times}$. Then the constancy of $W_{0}^{\kappa}(t)$, $|t-t^{*}|<{\varepsilon}$, follows from [15, Lemmas 13 and 15], provided, we establish the following transversality condition. Choose a sufficiently large integer $s$. For each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$, we set $I_{z}=I^{cond}(C,z)/{\mathfrak{m}}_{C,z}^{s}\subset{\mathcal{O}}_{C,z}/{\mathfrak{m}}_{z}^{s}$, the quotient of the conductor ideal by the power of the maximal ideal, which can be viewed as the tangent cone to the stratum parameterizing equigeneric deformations (see [4, Theorem 4.15]). The condition we need reads
$h^{1}\left(C,{\mathcal{J}}_{Z/C}\otimes_{{\mathcal{O}}_{\widehat{C}}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}}\right)=0,$
where ${\mathcal{J}}_{Z/C}$ is the ideal sheaf of the zero-dimensional scheme $Z\subset C$ defined at each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)$ by the conductor ideal $I^{cond}(C,z)$ and at each point $w=w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}(t^{*})$ by the ideal
$I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ (\varphi\cdot C)_{w}\geq 2k_{i}^{\sigma}-1\}\subset{\mathcal{O}}_{C,w},$
The above $h^{1}$-vanishing lifts to ${\mathbb{P}}^{1}$ in the form
$h^{1}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}({\boldsymbol{d}}))=0,$ (15)
where
$\deg{\boldsymbol{d}}=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2)-(CK_{{\operatorname{Tor}}(P_{\Delta})}-n-1)$
$=n-1>-2=2g-2,$ (16)
and hence the required $h^{1}$-vanishing comes from the Riemann-Roch theorem.
(2) Suppose that $\xi\in\overline{\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ is as in Lemma 4.2(2i) with singularities of type $A_{2m}$ at some points $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}(t^{*})$, $m\geq k_{i}^{\sigma}$.
The germ at $\xi$ of ${\mathcal{M}}_{0,n}(\Delta)$ (defined by (3)) can be identified with the germ at 00 of $H^{0}(C,{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})$, which in turn can be naturally embedded into $\prod_{\sigma,i}\mathrm{Vers}(C,w_{i}^{\sigma})\times\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}\mathrm{Vers}(C,z)$. We claim that the image of the germ at 00 of $H^{0}(C,{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})$ in $\prod_{i,\sigma}\mathrm{Vers}(C,w_{i}^{\sigma})\times\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}\mathrm{Vers}(C,z)$ intersects transversally with $\prod_{\sigma,i}EG_{{\operatorname{Tor}}(\sigma)}(C,w_{i}^{\sigma})\times\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}EG(C,z)$. Hence, this intersection ${\mathcal{V}}$ is smooth of the expected dimension $n$ since the latter product is smooth (see Section 2.3, items (c) and (d)). The required transversality amounts to the validation of the relation
$h^{1}(C,{\mathcal{J}}_{Z_{1}/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})=0,$ (17)
where the zero-dimensional scheme $Z_{1}\subset C$ supported at $\widehat{\boldsymbol{w}}(t^{*})\cup{\operatorname{Sing}}\hskip 1.70709pt(C)$ and defined by the ideals $I^{cond}(C,z)$ at all $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$ and by the ideals
$\{\varphi\in{\mathcal{O}}_{C,w_{i}^{\sigma}}\ :\ {\operatorname{ord}}\varphi\big{|}_{C,w_{i}^{\sigma}}\geq 2k_{i}^{\sigma}-1+2\delta(C,w_{i}^{\sigma})\}\quad\text{for all}\ w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}(t^{*})$ (18)
(cf. Section 2.3, items (c) and (d)). Relation (17) is equivalent to
$h^{1}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}({\boldsymbol{d}}_{1}))=0,$
where $\deg{\boldsymbol{d}}_{1}$ is given by (16), and hence it holds by the Riemann-Roch theorem.
Note that $\xi$ is chosen as a generic member of the wall, that is, it represents a smooth point of the wall. Let us show that the germ ${\mathcal{V}}_{1}$ of the wall at $\xi$ has codimension one in ${\mathcal{V}}$. Pick a point $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}(t^{*})$, where ${\Phi}(\widehat{C})$ has a singularity of type $A_{2m}$, $m\geq k_{i}^{\sigma}$, and consider the intersection ${\mathcal{V}}_{2}$ of the image of the germ at 00 of $H^{0}(C,{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})$ in $\prod_{\sigma^{\prime},i^{\prime}}\mathrm{Vers}(C,w_{i^{\prime}}^{\sigma^{\prime}})\times\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}\mathrm{Vers}(C,z)$ with
$\prod_{(\sigma^{\prime},i^{\prime})\neq(\sigma,i)}ES_{{\operatorname{Tor}}(\sigma^{\prime})}(C,w_{i^{\prime}}^{\sigma^{\prime}})\times ES^{fix}_{{\operatorname{Tor}}(\sigma)}(C,w_{i}^{\sigma})\times\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}EG(C,z).$ (19)
It has dimension $\dim{\mathcal{V}}_{2}\geq\dim{\mathcal{V}}_{1}-1$. We claim that $\dim{\mathcal{V}}_{2}\leq\dim{\mathcal{V}}-2=n-2$. Indeed,
$h^{1}(C,{\mathcal{J}}_{Z_{2}/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})=0,$ (20)
where the zero-dimensional scheme $Z_{2}$ supported at $\widehat{\boldsymbol{w}}(t^{*})\cup{\operatorname{Sing}}\hskip 1.70709pt(C)$ coincides with $Z_{1}$ (appearing in (17)) along $(\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}\})\cup{\operatorname{Sing}}\hskip 1.70709pt(C)$ and is defined at $w_{i}^{\sigma}$ by the ideal
$\{\varphi\in{\mathcal{O}}_{C,w_{i}^{\sigma}}\ :\ {\operatorname{ord}}\varphi\big{|}_{C,w_{i}^{\sigma}}\geq 2k_{i}^{\sigma}+1+2\delta(C,w_{i}^{\sigma})\}$
(cf. Section 2.3, item (b)). To see this, as in the preceding paragraph we rewrite (20) in the form
$h^{1}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}({\boldsymbol{d}}_{2}))=0,$
where
$\deg{\boldsymbol{d}}_{2}=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2)-(CK_{{\operatorname{Tor}}(P_{\Delta})}-n)-1=n-3>-2.$
Thus, by the Riemann-Roch theorem, we have
$\dim{\mathcal{V}}_{2}\leq h^{0}(C,{\mathcal{J}}_{Z_{2}/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})=(n-3)+1=n-2,$
as required. Furthermore, relation (20) yields the existence of the germ of a smooth path through 00 in ${\mathcal{V}}$ transversally intersecting ${\mathcal{V}}_{1}$ and regularly projecting onto a smooth path through $\widehat{\boldsymbol{w}}(t^{*})$ in $\operatorname{Men}(\Delta)$.